# Life Sequence Transformer: Generative Modelling for Counterfactual Simulation

### v1

@@ 5 Empirical Analysis
In this section, we present the next-token prediction task metrics and evaluate the model’s ability to forecast out-of-sample for years beyond the training period. We also evaluate performance through a series of causal inference benchmarking exercises by comparing the generated life sequences against well-established empirical findings from labour economics. All experiments were run on both model variants, i.e. the encoder-decoder and the decoder-only architectures, and on the test subset of our data. In the main text, we present the best results and provide a discussion of how the two models compare across tasks, highlighting the strengths and weaknesses of each approach. The remaining results are included in Appendix E.
@@ 5.1 Next-Token Prediction Metrics
We begin by presenting the model’s performance on the next-token prediction task over the test subset. This task measures the model’s ability to correctly generate the next element in a life sequence, conditioned on all previous tokens and the encoded background information. Training each model took approximately 9 hours per epoch for the training phase and 1.5 hours per epoch for validation, totalling 10.5 hours per epoch. Since we train models between 15 and 20 epochs, depending on convergence, the total training time ranges between 158 and 210 hours per model.
Table 1 reports a standard set of classification metrics computed across the entire sequence dimension and vocabulary, excluding padding tokens. Each row corresponds to a different amount of observed history provided as input. Specifically, we generate the continuation of a life sequence after observing either zero, one, five, or ten years of past events. The metrics include macro-averaged accuracy, precision, recall, and F1 score, each evaluated using top 1 predictions, i.e. only the highest probability considered to find the correct label. We also report the square root cross-entropy loss, i.e. perplexity, a widely used measure in language modelling.
Both model variants achieve similar results on these metrics, so we report only the encoder-decoder results for brevity. As expected, the more historical context the model receives, the better it can predict future tokens. This is consistent with the intuition that individuals often follow relatively stable life patterns, e.g. remain within certain industries over time. Access to a longer sequence of past events enables the model to learn person-specific trajectories and reduces uncertainty about what is likely to happen next.
Table 1: Encoder-Decoder test set metrics for next-token prediction, macro-averaged across the vocabulary, conditioning on 0, 5, 1 and 10 initial years of sequence.
Accuracy Precision Recall F1 Score Perplexity
0 known years 0.6236 0.6874 0.6236 0.6445 0.725
1 known year 0.6763 0.7179 0.6763 0.6872 0.696
5 known years 0.6869 0.7251 0.6869 0.6967 0.675
10 known years 0.7023 0.7382 0.7023 0.7108 0.6489
@@ 5.2 Causal Benchmarks
We evaluate the model’s capacity for generating causal relationships by testing whether it reproduces well-established causal effects from the labour economics literature and national policy. All generation experiments are run on batches of 8 sequences. On a GPU, generating the next token takes approximately 0.2 seconds per batch, while on a CPU it takes about 2 seconds. We design three empirical exercises (more details are provided in Appendix D):
The first exercise is inspired by the finding that birth month causally impacts retirement timing (Ardito et al., 2020). The mechanism behind this is that institutional timing (e.g., school starting in September, job transitions peaking in January) causes individuals born earlier in the year to begin and end schooling later, delaying their entry into the labour market. This delay propagates throughout life and leads to a slightly later retirement, observable in the timing of pension claims. We assess whether the model replicates this subtle but well-documented shift in retirement age based on birth month.
Figure 2 compares real and simulated retirement ages across birth cohorts, distinguishing between individuals born in January and December. The model successfully captures the expected empirical pattern: individuals born in December tend to retire earlier than those born in January, reflecting institutional rules that influence retirement timing. While it slightly overestimates the average retirement age, these predictions remain well within the confidence intervals of the real data, indicating that the model aligns closely with observed outcomes.
Figure 8 shows that the decoder-only model overestimates the mean of retirement ages more, while still preserving the expected negative effect of birth cohort on retirement age. Its weaker predictive accuracy highlights a consistent pattern across exercises where background information plays a central role: the decoder-only model, which receives background variables only as part of the initial sequence, struggles with their long-term influence. The encoder-decoder model incorporates this information via cross-attention, allowing for more targeted attention to the time-invariant attributes of the individual, at the cost of allocating fewer parameters to self-attention within the sequence.
Figure 2: Real and encoder-decoder model-generated (blue) average pension ages—measured in months—for male individuals born between 1940 and 1950, stratified by birth month (January = M1, December = M12). Each point indicates the cohort-specific average retirement age for individuals born in January or December, with vertical bars representing a one standard deviation range.
The second exercise relates to the motherhood penalty, i.e. the long-term career setbacks women face after having children, often including reduced earnings and detachment from the labour market (Casarico and Lattanzio, 2023; Kleven et al., 2019). We simulate mother’s careers from the moment of the first maternity leave event and compare the post-child trajectory with observed empirical trends.
Figure 3 compares the real and simulated income trajectories of mothers following their first childbirth. Synthetic sequences are generated conditionally on the sequence up to the maternity leave year. The model successfully captures the key features of the observed trends, closely replicating the average post-birth income dynamics. For this task, we report results from the decoder-only model, which demonstrates superior performance. We attribute this to the fact that background variables, like area of birth or age, play a limited role in explaining the after-birth trend, making the task particularly well-suited for models focused on local sequence patterns rather than long-range context. In contrast, Figure 9 shows how the encoder-decoder model’s predictive accuracy gradually declines the further we move away from the observed sequence.
Figure 3: Real and decoder-only model-generated trajectories of average monthly income for working mothers in a 10-year window before and after their first maternity leave episode (Years From Maternity = 0). The solid lines represent the average monthly income, while the shaded areas denote the one standard deviation range.
@@
Lastly, we evaluate the ability to generate the impact of an unemployment benefit policy rule in Italy. Until 2017, Italy’s mobility allowance granted income support to workers displaced due to firm closure. The benefit duration had discontinuities based on age (cutoff at 40 and 49) and region (South vs. North/Central), creating sharp jumps in benefit duration entitlement and reception (Brunello and Miniaci, 1997).
@@
Figure 4 compares the real and simulated average duration of mobility allowance receipt. In the real data, we observe a statistically significant jump at the age 40 threshold, reflecting the policy rule. The simulated data captures the overall upward trend in benefit duration with age, suggesting that the model has learned the general relationship between age and benefit duration. However, it falls short of reproducing the sharp discontinuity observed at the age 40 threshold in the real data. While the simulated sequences reflect the correct direction of the trend, they smooth over the abrupt policy-induced change, indicating a limitation in the model’s sensitivity to this rule-based cutoff.
This underperformance is likely due to the low frequency of mobility allowance events in the training data, i.e. less than 1% of all sequences (see Figure 5), which limits the model’s ability to learn the underlying policy pattern. The decoder-only model, shown in Figure 12, fails to produce any significant change in benefit duration at the 40-year cutoff. This again highlights the limitations of the decoder-only approach in tasks where access to and emphasis on background variables, like age and region, are crucial for capturing policy-driven effects.
@@
Figure 4: Real and encoder-decoder model simulated average duration with 90% confidence interval of mobility allowance receipt duration as a function of worker age at job loss in months in a 10-year bandwidth around the 40 years old threshold
@@ 5.3 Out-of-sample Generation
We evaluate the model’s out-of-sample performance by investigating the coherence of generated life sequences beyond the temporal window of the training data, i.e. after 2015. In this setting, the model is conditioned on observed sequences up to 2015 and generates tokens into the future. Though not explicitly necessary for counterfactual simulation, good out-of-sample performance indicates robustness, a desirable property for a model expected to simulate plausible scenarios.
@@
Given the absence of ground truth data for these later years, our evaluation focuses on syntactic and structural correctness rather than predictive accuracy. Specifically, we examine the generated sequences for internal consistency, valid token ordering, and the absence of structural violations—such as overlapping events, misaligned episode boundaries, or implausible transitions. This assessment allows us to verify that the model preserves the learned "grammar" of life sequences when extrapolating beyond the observed data.
@@
Figure 13 presents the density of years in which the first structural failure occurs for the decoder-only model, considering only sequences where a failure did occur within a 20-year generation window. The results show that most decoder-only sequences begin to fail around year 7 or 8, though a non-negligible share persists further, with successes becoming exponentially less likely as the generation horizon extends, and 5% of sequences remain structurally valid after 20 years. In contrast, the encoder–decoder model (Figure 14) exhibits earlier degradation, with most failures concentrated around year 4, very few sequences remaining valid beyond year 10 and none after year 20. These patterns confirm that decoder-only architectures are more robust for long-range generation tasks, maintaining structural coherence significantly longer than encoder–decoder counterparts.
6 Conclusion
We propose a novel encoding method for converting longitudinal administrative data into sequences and train two generative life sequence transformer models to simulate coherent and realistic counterfactual trajectories. The models are trained on a large-scale administrative dataset and optimised for next token prediction. Despite their relatively small size, the models show promising results in replicating well-known causal findings, showing the potential of lightweight architectures in this domain. We highlight the strengths of two architectural paradigms—decoder-only and encoder–decoder—across a range of causal generative tasks. In an out-of-sample sequence generation setting, the model remains functional, although token prediction accuracy and sequence coherence degrade increasingly with temporal distance from the training data.
Despite its promising performance, our model presents some limitations. As shown in the mobility allowance duration exercise, the model has less accurate generative capacities when used to generate sequences with events that are rare in the training data. This limitation is particularly relevant given the non-open nature of administrative data, which limits dataset sizes. This highlights the importance of fine-tuning the model for specific subpopulations or tasks. Additionally, the quality of sequence generation is sensitive to model architecture, introducing variability in counterfactual simulation tasks. Careful model selection and design are therefore essential; researchers must evaluate the trade-offs between architectures and systematically test and compare them to ensure robust results.
While the model offers a powerful tool for simulating policy scenarios and exploring counterfactual life trajectories, it raises important ethical and societal concerns. Predictive modelling of individual life courses—especially in domains such as employment, healthcare, or social services—carries the risk of reinforcing structural biases in the training data. There is also the potential for misuse, for example, in automated decision-making systems that influence resource allocation or eligibility for benefits, without appropriate human oversight or fairness safeguards. Out-of-sample generation into the future introduces further moral and philosophical concerns: guiding individuals’ lives based on generically generated futures could contribute to normative pressures or reductionist narratives. Our model’s limitations in long-range future prediction constrain such misuse, as it performs best when simulating counterfactuals within known historical periods. Nonetheless, any application of generative models to human trajectories must be governed by strict considerations of fairness, privacy, and security. Addressing these challenges will be critical to ensuring that such models are used responsibly, transparently, and in ways that support robust causal inference.
@@ Appendix A Difference-in-differences
Difference-in-differences (DiD) is one of the more commonly used causal identification strategies to evaluate the effects of policy interventions by comparing changes in an outcome variable between a treatment group and a control group, both before and after the intervention. Here, the main assumption is that the treated and control groups would have followed parallel trends over time in the absence of the policy or event under evaluation. This assumption, while critical, is rarely validated and often debated in empirical applications.
Card and Krueger [1994] on minimum wage in New Jersey is considered one of the most famous DiD studies; the authors compared employment in the fast food sector in New Jersey and in Pennsylvania, in February 1992 and in November 1992, after New Jersey’s minimum wage rose from \$4.25 to \$5.05 in April 1992. Observing a change in employment in New Jersey only, before and after the treatment, would fail to control for omitted variables such as weather and macroeconomic conditions of the region. By including Pennsylvania as a control in a difference-in-differences model, any bias caused by shared variables between New Jersey and Pennsylvania is implicitly controlled for, even when these variables are unobserved. Assuming that New Jersey and Pennsylvania have parallel trends over time, Pennsylvania’s change in employment can be interpreted as the change New Jersey would have experienced had it not increased the minimum wage, and vice versa. The evidence suggested that the increased minimum wage did not induce a decrease in employment in New Jersey, contrary to what some economic theory would suggest.

### v2

@@ 6 EMPIRICAL ANALYSIS
In this section, we present the next-token prediction task metrics and evaluate the model’s ability to forecast out-of-sample for years beyond the training period. We then evaluate performance through a series of causal inference benchmarking exercises by comparing the generated life sequences against well-established empirical causal findings from labour economics. All experiments are run on the test subset of our data and code can be found in github/life-sequence-transformer.
@@ 6.1 Next-Token Prediction Metrics
We begin by evaluating the model on a next-token prediction task over the test subset, which measures how accurately it generates the next element in a life sequence given prior tokens and background information. Training each model took approximately 9 hours per epoch for the training phase and 1.5 hours per epoch for validation, totalling 10.5 hours per epoch. The final model was trained for 20 epochs for a total training time of 210 hours on a single NVIDIA RTX A4000 GPU with 16GB of VRAM.
Table 1 reports a standard set of classification metrics computed across the entire sequence dimension and vocabulary, excluding padding tokens. Each row corresponds to a different length of observed history provided as input. Specifically, we generate the continuation of a life sequence after observing either zero, one, five, or ten years of past events. The metrics include macro-averaged accuracy, precision, recall, and F1 score, each evaluated using top-1 predictions, i.e., consider only the highest probability to assign the predicted token. We also report the square root cross-entropy loss, i.e., perplexity, a widely used measure in language modelling.
As expected, the more historical context the model observes, the better it can predict future tokens. This is consistent with the intuition that individuals often follow relatively stable life patterns after a certain age, e.g., remain within certain industries over time. Access to a longer sequence of past events enables the model to learn person-specific trajectories and reduces uncertainty about what is likely to happen next.
@@ 6.2 Causal Benchmarks
We evaluate the model’s capacity for generating causal relationships by testing whether it reproduces well-established effects from the labour economics literature and national policy. To do so, we truncate observed individual life sequences at specific cutoff years and condition the model on this truncated history. The model is then tasked with generating the continuation of the sequence forward in time. We only truncate at the level of whole years, so that the conditioning context always consists of a realistic, complete history up to the cutoff point. In each experiment, we choose the cutoff year to coincide with the onset of the event of interest (i.e., unemployment, maternity, or retirement), so that the subsequently generated trajectory can be compared to known empirical effects. Additional details of this procedure are provided in Appendix D. All generation experiments are run on batches of 8 sequences. On a GPU, generating the next token takes approximately 0.2 seconds per batch, while on a CPU it takes about 2 seconds.
$\text{ATE}_{\text{emp}}$ $\text{ATE}_{\text{model}}$ $\Delta$ $n$
Pension (1y cutoff) 7.47 [2.50, 12.23] 9.10 [4.95, 13.26] 1.62 [-0.36, 3.67] 1165
Pension (4y cutoff) 7.47 [2.42, 12.59] 11.02 [7.07, 15.14] 3.55 [ 2.09, 5.07] 1165
Unemployment $\pm$12m 4.18 [1.46, 6.67] 0.21 [-2.17, 2.55] -3.98 [-6.38, -1.66] 153
Unemployment $\pm$48m 3.41 [1.76, 4.90] 2.49 [1.08, 3.89] -0.92 [-2.18, 0.54] 556
Unemployment $\pm$96m 7.25 [5.70, 8.70] 7.09 [5.86, 8.39] -0.16 [-1.36, 0.97] 1168
Unemployment $\pm$144m 8.27 [7.02, 9.51] 8.92 [7.76, 10.02] 0.65 [-0.34, 1.67] 1502
Maternity 0–3 years (%) -25.1 [-28.3, -22.2] -22.6 [-24.6, -20.5] 2.45 [-1.20, 6.20] 631
Maternity 0–5 years (%) -19.1 [-22.0, -16.5] -18.8 [-20.7, -16.8] 0.31 [-2.90, 3.80] 631
Maternity 0–10 years (%) -11.0 [-13.7, -8.30] -13.4 [-15.6, -11.2] -2.35 [-5.90, 1.10] 631
Table 2: Empirical and model-based average treatment effects (ATEs) point estimates with 95% confidence interval in brackets. $\Delta$ is the signed difference between model and empirical ATEs. $n$ denotes the number of individuals in the analysis sample. For details on how ATEs are constructed in each experiment, see Appendix D.
@@
First, we evaluate whether the model captures the impact of an unemployment benefit policy rule in Italy. Until 2017, Italy’s mobility allowance granted income support to workers displaced due to firm closure, with benefit duration varying discontinuously by age (cutoffs at 40 and 49) and by region (South vs. North/Central) (Brunello and Miniaci, 1997). For each individual, we truncate the sequence at the year when displacement is first observed and provide this truncated history to the model. The model must then generate forward, including the timing and duration of subsequent unemployment and benefit receipt.
@@
Figure 2 shows the simulated average duration of mobility allowance receipt as a function of age at displacement. The simulated data captures the overall upward trend in benefit duration with age, suggesting that the model has learned the general relationship between age and benefit duration. However, it falls short of reproducing the sharp discontinuity observed at the threshold, instead smoothing over the abrupt policy-induced change. One likely reason is data scarcity: fewer than 1% of sequences involve a mobility allowance spell (see Figure 6), which restricts the amount of policy-relevant signal available for training.
@@
Figure 2: Model-generated average duration with 90% confidence interval of mobility allowance receipt duration as a function of worker age at job loss in months in a 10-year bandwidth around the 40-year-old threshold.
A more formal discontinuity analysis confirms this pattern. Table 2 reports local average treatment effects (ATEs) for different bandwidths around the 40-year cutoff. At narrow bandwidths (e.g., $\pm 12$ months), the model ATE fails to significantly capture the discontinuity. As the window expands to $\pm 48$ and $\pm 96$ months, however, the model’s ATEs move closer to the empirical benchmark, with discrepancies shrinking to below one month and confidence intervals overlapping zero. At the widest bandwidth ($\pm 144$ months), both the empirical and model estimates are around 8–9 months, with a modest and statistically insignificant difference.
The second exercise examines the motherhood penalty, i.e., the long-term career setbacks women face after having children, often including reduced earnings and labour market detachment (Casarico and Lattanzio, 2023; Kleven et al., 2019). Here, we simulate women’s careers from the moment of first maternity. Specifically, we truncate sequences either (i) at the year of the first observed maternity leave (offset–0), or (ii) one year before that event (offset–1). The latter setup allows maternity leave itself to be generated or not, and enables counterfactual analysis by comparing career paths with and without maternity.
Figure 3 compares the real and simulated income trajectories of mothers following their first childbirth (see a breakdown by sector in Figure 12). Synthetic sequences are generated conditional on the real sequence up to the maternity leave year. The model successfully captures the key features of the observed trends, closely replicating the average post-birth income dynamics.
Figure 3: Real and model-generated trajectories of average annual earnings for working mothers in a 10-year window before and after their first maternity leave episode (offset–0). The solid lines represent average annual earnings, while the shaded areas denote the one standard deviation range.
To estimate the treatment effect of childbirth on earnings, we use an event-study design with controls for age and calendar year, as in Kleven et al. (2019). Because fertility is endogenous, we construct counterfactual trajectories using both propensity-score matched non-mothers and model-generated no-maternity paths (offset–1 simulations). The details of this methodology are presented in Appendix D.2. These event-study coefficients are then summarised into average percentage income penalties over post-birth horizons, making them directly comparable to the ATEs reported for pension and unemployment.
The results show that the model reproduces the empirical child penalties closely across all horizons (Table 2). What is especially interesting is the pattern of differences: the model tends to understate the immediate short-run penalty but generates a more persistent effect in the medium and long run. In other words, the simulated trajectories suggest a smaller initial income drop but slower recovery, implying that the penalty lingers for longer.
Figure 10 illustrates how the model-generated simulated controls (non-maternity paths) compare to real mothers’ earnings trajectories, with propensity score–matched mothers included as a comparison. The simulated controls track the counterfactual trend, reinforcing the finding that the model not only captures the observed income losses of mothers but also produces realistic no-maternity baselines.
The final exercise investigates whether the model replicates the well-documented effect of birth month on retirement timing (Ardito et al., 2020). Institutional features, such as school starting dates and labour market entry, generate systematic differences in retirement timing between individuals born in January versus those born in December. To test this, we truncate observed sequences prior to retirement. In one variant, we cut off at the year immediately before the observed retirement (so the model must decide whether retirement occurs in the following year or later). In another variant, we truncate four years before retirement, which increases the uncertainty in retirement timing and allows us to compare under different predictive horizons.
Figure 4 shows results with the one-year cutoff. Here, the model reproduces the negative effect of December birth on retirement age, indicating robustness of the signal. However, a clear upward bias emerges. This suggests that conditioning the model too close to the event may shift its internal expectations and effectively push retirement further into the future.
Figure 4: Real (circle dots) and model-generated (triangle dots) average retirement ages, measured in months, for male individuals born between 1940 and 1950, using a cutoff one year before the observed retirement year. Each point indicates the cohort-specific average retirement age for January (M1) and December (M12), with vertical bars representing one standard deviation.
Figure 5 reports results with the four-year cutoff. The model captures the expected empirical pattern: December-born individuals retire earlier than January-born peers, and the magnitude of this birth-month effect is well preserved across cohorts. The upward bias in retirement ages is largely absent under this earlier cutoff, indicating that conditioning further from the retirement event leads to a closer alignment with the observed timing of pension entry.
Figure 5: Real (circle dots) and model-generated (triangle dots) average retirement ages, measured in months, for male individuals born between 1940 and 1950, using a cutoff four years before the observed retirement year. Each point indicates the cohort-specific average retirement age for January (M1) and December (M12), with vertical bars representing one standard deviation.
Table 2 formalises these findings. Both cutoffs generate significant differences in retirement age between treated (born in January) and untreated (born in December) individuals, consistent with the empirical benchmark. Yet the size of the model’s effect is sensitive to the conditioning horizon. Truncating the sequence further from the outcome offers a stronger test of counterfactual reasoning. Still, it tends to overstate effect sizes, illustrating how the length of the conditioning horizon shapes both predictive accuracy and the model’s interpretation of the causal nexus among life events.
@@ 6.3 Out-of-sample Generation
We evaluate the model’s out-of-sample performance by investigating the coherence of generated life sequences beyond the temporal window of the training data, i.e., after 2015. In this setting, the model is conditioned on observed sequences up to 2015 and generates tokens into the future. Good out-of-sample performance indicates robustness, a desirable property for a model expected to simulate plausible scenarios.
@@
Given the absence of ground truth data for these later years, our evaluation focuses on syntactic and structural correctness rather than predictive accuracy. Specifically, we examine the generated sequences for internal consistency, valid token ordering, and the absence of structural violations, such as overlapping events or misaligned episode boundaries (more details in Appendix D.4). This assessment allows us to verify that the model preserves the learned “grammar” of life sequences when extrapolating beyond the observed data.
@@
Figure 14 presents the density of out-of-sample generated years in which the first structural failure occurs, considering only sequences where a failure did occur within a 20-year generation window. The results show that sequence generation begins to fail around year 7 or 8, though a non-negligible share persists further, with successes becoming exponentially less likely as the generation horizon extends, and 5% of sequences remain structurally valid after 20 years.
7 DISCUSSION
We propose a novel encoding method for converting longitudinal administrative data into sequences and train a generative life sequence Transformer model. Despite its relatively small size, the model learns to reproduce realistic and causally consistent labour market histories and generates coherent synthetic life paths conditional on past trajectories, showing the potential of Transformer architectures in this domain.
Through empirical exercises, we study the strengths and limitations of our model. On the one hand, it successfully reproduces causal patterns, from birth-month effects on retirement to the motherhood penalty. On the other hand, the exercises reveal systematic sensitivities. Rule-driven discontinuities are smoothed over when the underlying events are rare in the data; treatment effects vary depending on how counterfactuals are defined, whether by comparing mothers to non-mothers or by truncating retirement histories at different horizons; and the timing of effects can shift, with the model sometimes understating short-run penalties but prolonging them over the longer term.
These findings highlight a broader challenge: counterfactual inference with generative models depends on how we define treatment, select controls, and decide where to cut the input history. These are not just technical choices but conceptual ones, raising fundamental questions about what constitutes a life sequence counterfactual. They also suggest future research on developing principled methods for introducing researcher-imposed variation into simulated life sequences, ensuring theory-guided design of counterfactual scenarios.
Finally, while our results show encouraging alignment with empirical findings, they also remind us that predictive modelling of individual life courses carries ethical and societal risks. Structural biases present in training data will be reinforced. Without appropriate human oversight, there is also the potential for misuse, for example, in automated decision-making systems that influence resource allocation or eligibility for benefits. Future applications must combine methodological advances with safeguards for fairness, privacy, and responsible use.
Acknowledgements
The authors acknowledge European NextGenerationEU funding through the PNRR project “Age-It - Ageing well in an ageing society”, code PE0000015, MUR 341/2022 tender.
Supplementary Materials
Counterfactual inference is a cornerstone of scientific reasoning in disciplines such as political science Sekhon (2009), economics (Heckman, 2008), history (Levy, 2015), and psychology (Foster, 2010), and has also sparked ongoing methodological debate (Cartwright, 2007a). Counterfactual identification requires constructing hypothetical scenarios that resemble what would have happened in the absence of an event assumed to be causal (Rubin, 2005). Such a scenario allows quantifying the impacts of specific events, like treatments or policies, or the underlying causes of observed outcomes. Typical research questions include: What is the causal effect of increasing the minimum wage on household consumption (Alonso, 2022)? What are the drivers of civil conflict (Acemoglu et al., 2020)?
However, randomisation rarely occurs naturally and can be challenging to implement even in controlled settings, due to ethical constraints and substantial costs (Thomas, 2016). As a result, economists and other researchers have developed strategies and mathematical models to identify suitable control groups without relying on random treatment assignment. Despite being widely adopted, these observational methods often rely on specific assumptions about the data and the underlying causal mechanisms, which are typically difficult to verify in practice.
@@ Appendix A Difference-in-differences
Difference-in-differences (DiD) is one of the most commonly used causal identification strategies for evaluating the effects of policy interventions by comparing changes in an outcome variable between a treatment group and a control group, both before and after the intervention. Here, the main assumption is that the treated and control groups would have followed parallel trends over time in the absence of the policy or event under evaluation. This assumption, while critical, is rarely validated and often debated in empirical applications.
Card and Krueger (1994) on minimum wage in New Jersey is considered one of the most famous DiD studies; the authors compared employment in the fast food sector in New Jersey and in Pennsylvania, in February 1992 and in November 1992, after New Jersey’s minimum wage rose from \$4.25 to \$5.05 in April 1992. Observing a change in employment in New Jersey only, before and after the treatment, would fail to control for omitted variables such as weather and macroeconomic conditions of the region. By including Pennsylvania as a control in a difference-in-differences model, any bias caused by shared variables between New Jersey and Pennsylvania is implicitly controlled for, even when these variables are unobserved. Assuming that New Jersey and Pennsylvania have parallel trends over time, Pennsylvania’s change in employment can be interpreted as the change New Jersey would have experienced had it not increased the minimum wage, and vice versa. The evidence suggested that the increased minimum wage did not lead to a decrease in employment in New Jersey, contrary to what some economic theories would suggest.